# On the Scarcity of Dense Cores ( cm) in High Latitude Planck Galactic Cold Clumps

### 26 Jan 2024 v1

@@ Abstract
High-latitude () molecular clouds have virial parameters that exceed 1, but whether these clouds can form stars has not been studied systematically. Using JCMT SCUBA-2 archival data, we surveyed 70 fields that target high-latitude Planck galactic cold clumps (HLPCs) to find dense cores with density of – cm and size of pc. The sample benefits from both the representativeness of the parent sample and covering densest clumps at the high column density end ( cm). At an average rms of 15 , we detected Galactic dense cores in only one field, G6.04+36.77 (L183), while also identifying 12 extragalactic objects and two young stellar objects. Compared to the low-latitude clumps, dense cores are scarce in HLPCs. With synthetic observations, the densities of cores are constrained to be cm, should they exist in HLPCs. Low-latitude clumps, Taurus clumps, and HLPCs form a sequence where a higher virial parameter corresponds to a lower dense core detection rate. If HLPCs were affected by the Local Bubble, the scarcity should favor turbulence-inhibited rather than supernova-driven star formation. Studies of the formation mechanism of the L183 molecular cloud are warranted.
=30
@@ 1 Introduction
The high latitude (HL hereafter) of the Milky Way, also called the “Underwater Iceberg” guards its secrets about molecular gas and star formation, due in part to the limited scope of previous CO surveys (Xu et al., 2021). Observational challenges essentially originate from the large area of the HL, leading to much longer integration times compared to those required for blind surveys of the Galactic plane. The Planck satellite provides an unprecedented all-sky census of the coldest (6–20 K with a median value of 14 K; Planck Collaboration et al., 2011a) Galactic objects by combining the highest frequency channels of the Planck survey 353–857 GHz (i.e., 850–350 m) with the far-infrared IRAS 100 m data (Neugebauer et al., 1984; Miville-Deschênes & Lagache, 2005). As a result, the Planck team has cataloged 13,188 Planck Galactic Cold Clumps (PGCCs; Planck Collaboration et al., 2016), including 793 with absolute value of Galactic latitude higher than 30 (HLPCs hereafter). Benefiting from the unbiased nature of its parent sample, the 793 HLPCs are the least biased sample of HL cold dust clumps, therefore serving as a foundation for studying the properties of the HL molecular gas and investigating the initial condition of star formation (Wu et al., 2012; Liu et al., 2013).
@@
Our previous work performed a CO/CO/CO (1-0) survey towards 41 Early Cold Cores (ECCs; i.e., most reliable detections of PGCCs with signal-to-noise ratio ; Planck Collaboration et al., 2011b) with the PMO 13.7 m millimeter-wavelength telescope (Xu et al., 2021). Although detected CO cores have a typical density of several times cm, consistent with what has been found in nearby molecular cloud cores (Myers et al., 1983; Myers & Benson, 1983; Myers, 1983; Benson & Myers, 1983, 1989), the turbulent energy is significantly higher than the gravitational energy, with median virial parameters of (Xu et al., 2021). Therefore, our CO surveys unveiled a highly turbulent, diffuse molecular gas environment as a first glimpse of the initial conditions of star formation in the HL clouds.
@@
Stars form in dense cores (Shu et al., 1987) with a typical size of pc and density of – cm (Ward-Thompson et al., 1994, 1999; Kirk et al., 2005). The CO (1-0) transitions suffer from optical thickness and depletion at low temperature, so it is hard to probe the densest regions of molecular clouds. For example, a high-latitude cloud L1780 shows a cometary morphology and a CO core (Toth et al., 1995), but contains no dense core in our survey (field G358.96+36.81). Furthermore, the angular resolution of Planck used for the extraction of PGCCs is ′, corresponding to 0.3 pc at a typical distance of HL clouds of 200 pc (Xu et al., 2021), which is marginal for resolving dense cores. Working at 850 m with an effective beam FWHM of 146, Submillimetre Common-User Bolometer Array 2 (SCUBA-2; Dempsey et al., 2013) provides times better resolution than the Planck, pinpointing cold dense cores inside the molecular clouds.
@@ 2.1 Sample Selection
A thorough search of the SCUBA-2 850 m data in the JCMT Science Archive111https://www.cadc-ccda.hia-iha.nrc-cnrc.gc.ca/en/jcmt/, cross-matching with 793 HLPCs that satisfy the latitude criterion of , gave 138 observing fields in total. We dismiss six of the fields which have limited integration time or non-standard scan modes. Different scan patterns includes Constant Velocity Daisy patterns (CV Daisy) and Rotating Curvy Pong patterns (Curvy Pong), so the field offsets vary significantly between different patterns. We make sure that the observing fields cover the peak of PGCCs at 353 GHz. We also check superposition or repetition: if two fields cover the same PGCC, we choose the one with the higher sensitivity. After the work flow, a total of 70 SCUBA-2 fields are selected as the sample in this work.
@@ 2.2 Sample Properties
70 HLPCs are shown with green crosses, overlaid on the background Planck 353 GHz (850 m) continuum emission in Figure 1. White rectangles outline the CO emission regions defined by Dame et al. (1987). The clump-averaged H column density was calculated assuming a dust emissivity model at 857 GHz (Planck Collaboration et al., 2016). The distributions of three samples, all the PGCCs, 793 high-latitude PGCCs, and 70 HLPCs are plotted as gray, blue and orange histograms in Figure 2, respectively. The span of the 70 HLPCs has been evenly sampled, ensuring a similar distribution with its parent sample, 793 HL PGCCs. More importantly, the studied sample includes the densest clumps at the high-density end ( cm). Considering that the denser clumps should be more likely to produce dense cores, we’ve covered the complete HLPCs where dense cores could form.
@@ 2.3 Distance Estimation
Distance is always a difficult quantity to estimate in astronomy. Previous studies have estimated the distances of HL molecular clouds to be 100 pc from the velocity dispersion and the scale height of an ensemble of clouds (Blitz et al., 1984; Magnani et al., 1985). The star counting confirms that the HL molecular clouds are indeed nearby objects with upper limit ranges from 125 to 275 pc (Magnani & de Vries, 1986). Using Strömgren photometry, Franco (1989) derive the distances of several HGaL clouds to be 100–230 pc. Both the small and the lack of a double-sine wave signature in the distribution on the (Galactic longitude ) plane demonstrate that HL molecular gas belongs to the local ISM and is too close to the Sun for Galactic rotation to modulate the velocities (Magnani et al., 1996).
Recently, the Gaia satellite has provided new photometric measurements towards galactic stars (Gaia Collaboration et al., 2016). Together with 2MASS and Pan-STARRS 1 optical and near-infrared photometry, Gaia DR2 parallaxes can help to infer distances and reddenings of 800 million stars. These stars trace the reddening on a small patch of the sky, along different lines of sight and different distance intervals, allowing us to build a 3D dust reddening map (Green et al., 2014, 2019). In a given direction, a jump of dust reddening is expected at a distance where there is a dust clump. Distances are estimated by this method and are listed in column (9) of Table A1, with an average value of pc, indicating that HLPCs are mostly local ISM. Adopting the Eq. 1 in Xu et al. (2021) and considering that the Sun is 10 pc above the Galactic midplane (Griv et al., 2021), the altitude from the midplane (in unit of pc) is calculated from where is the distance and is the latitude, and listed in column (10). We note that some fields may contain extragalactic objects, so the distance should be only for foreground Galactic dust.
@@ 3.1 Source Extraction
We adopt the dendrogram algorithm (Rosolowsky et al., 2008) to extract dense structures, and then measure their integrated flux, peak flux, size and position by 2D Gaussian fitting. The details of the algorithm parameter settings and the source extraction procedure are introduced in Appendix B. Within the 70 input fields, we have initially detected a total of 20 sources that belong to 15 SCUBA-2 observing fields. The field names and extracted sources are listed in columns (1)–(2) of Table 1. Central coordinates, standard deviation of the deconvolved major and minor axes, position angle, integrated flux, and peak intensity are listed in columns (3)–(8).
@@ 3.2 Source Identification and Dense Core Definition
The CV Daisy mode of observation can produce artifacts (Liu et al., 2018; Eden et al., 2019) which are extracted by the algorithm as false source detections. Therefore, we additionally require that both the peak intensity after being smoothed to the Planck beam and the total flux of the source at 353 GHz, are lower than those of the parent PGCC. This results in two sources, G50.41-35.40 SMM1 and G197.98+33.10 SMM1, being classified as artifacts and excluded from further analyses. We note that the flux given by the PGCC catalog should be from the cold residual map (Planck Collaboration et al., 2016), so the original Planck flux at 353 GHz could be even larger. However, considering the cold nature of dense cores, this should contribute little to the warm component.
We cross-match the true detections within 1′ using SIMBAD222http://simbad.u-strasbg.fr/simbad/. We find that only one field, G6.04+36.77, pointing towards the molecular cloud L183, contains three resolved (or marginally resolved) sources, which were previously identified as low-mass prestellar cores (Dickens et al., 2000; Pagani et al., 2003). Our other detections are unresolved as point sources, classified as either young stellar object (YSO) or extragalactic object (point sources) - including gravitational lensed galaxy (LeG), BL Lacertae object (BLL), protocluster of galaxies (PClG), active nuclei candidate (AGN), and quasar (QSO). The determined identifiers and references are given in column (9) of Table 1.
@@ 4.1 Dense Cores Are Scarce In HLPCs
Having only one detection among the 70 HLPCs highlights the scarcity of dense cores at these latitudes. To further confirm this discrepancy compared to the low-latitude () counterpart, 1235 observing fields from the JCMT Large Project “SCUBA-2 Continuum Observations of Pre-protostellar Evolution” (SCOPE; Liu et al., 2018; Eden et al., 2019) are used as a comparison group because of the following two reasons. 1) 21 HLPC fields in this work come from the SCOPE project so that they have been observed with comparable sensitivities; 2) the SCOPE observations serve as a representative sample of PGCCs, with similar distributions in distance, size, and temperature, and with complete column density coverage over cm (Liu et al., 2018).
Considering the beam dilution effects for marginally resolved sources, the column densities are corrected by a beam filling factor where is the Planck beam size at 857 GHz (Planck Collaboration et al., 2016) and is the intrinsic size deconvolved from the beam. For extended sources of which intrinsic size exceeds the beam size, is set to 1 and no correction is performed. As a result, the corrected column densities increase by a factor of 1.1 on average and 9 at maximum.
@@ 4.2 What Does “Dense” Mean?
The scarcity of detection does not necessarily indicate the scarcity of dense cores, primarily due to two factors: 1) the limited sensitivity of identifying sources and 2) the absence of large-scale flux in the SCUBA-2 data processing. Therefore, to understand what the scarcity of dense cores means, it is essential to clarify what “dense” means.
Prestellar cores are observed to have flat inner density gradients that approach beyond a few thousand au (Ward-Thompson et al., 1994, 1999; Kirk et al., 2005), which can be reproduced by a nonmagnetic and Plummer-like model (Whitworth & Ward-Thompson, 2001) as,
@@
Using the GBT+JVLA combined data, the temperatures of the three prestellar cores are observed to have a minor decrease toward the center of the core au (Lin et al., 2023). Therefore, we consider a constant temperature profile as K in the following discussion.
@@
In Figure 4, the background color map shows observed peak intensities across the 2-D parameter space of flat radius au and central density cm. The prestellar cores have been reported to have au (Ward-Thompson et al., 1999; Kirk et al., 2005), which are delineated by the left green dashed line in Figure 4. The right green dashed line marks 20,000 au, which corresponds to 0.1 pc.
@@
To further demonstrate this upper limit of density, we smooth the images to a resolution of 20″. With better sensitivity, two new cores (on in HLPC G159.41-34.37 and one in G161.87-35.76) are identified by the same algorithm and parameter input, which are called weakly detected cores. They have radii of about 6700 and 7500 au and averaged density of cm and cm, respectively. The two cores are labelled as orange rectangles in Figure 4. We also retrieve Herschel cold cores in L134 (HLPCs G4.13+35.75 and G4.17+36.67 in our survey), MBM12 (G159.21-34.28, G158.51-33.99, G159.14-33.79, G159.23-34.51, G159.41-34.37, G159.66-34.31), L1642 (G210.90-36.55), and LDN1780 (G358.96+36.81, G359.21+36.89) from Montillaud et al. (2015) which are labelled as green triangles. As they all lie in the gray region, below our sensitivity limit, these cold cores are not dense enough for detection and are consistently below the density limit of cm. As noted by Ward-Thompson et al. (2016), SCUBA-2 selects the densest cores from a population at a given temperature, which makes SCUBA-2 ideal for identifying those cores in Herschel catalogues that are closest to forming stars. So it is of great interest to study how these low-density cores form and whether they can still form stars or are transient objects.
As indicated by the black contours, peak intensity of a synthetic model is always below the corresponding outlined by white contours. The difference reflects the missing large-scale flux, which increases in importance from 0.06 to 0.31 dex with from 2500 to 20,000 au. Therefore, if we don’t consider missing flux, the density limit can be even lower, especially for those cores with larger flat radius.

### 22 Feb 2024 v2

# On the Scarcity of Dense Cores ( cm) in High-latitude Planck Galactic Cold Clumps

@@ Abstract
High-latitude () molecular clouds have virial parameters that exceed 1, but whether these clouds can form stars has not been studied systematically. Using JCMT SCUBA-2 archival data, we surveyed 70 fields that target high-latitude Planck Galactic cold clumps (HLPCs) to find dense cores with density of – cm and size of pc. The sample benefits from both the representativeness of the parent sample and its coverage of the densest clumps at the high column density end ( cm). At an average rms of 15 , we detected Galactic dense cores in only one field, G6.04+36.77 (L183) while also identifying 12 extragalactic objects and two young stellar objects. Compared to the low-latitude clumps, dense cores are scarce in HLPCs. With synthetic observations, the densities of cores are constrained to be cm should they exist in HLPCs. Low-latitude clumps, Taurus clumps, and HLPCs form a sequence where a higher virial parameter corresponds to a lower dense-core detection rate. If HLPCs were affected by the Local Bubble, the scarcity should favor turbulence-inhibited rather than supernova-driven star formation. Studies of the formation mechanism of the L183 molecular cloud are warranted.
=50
@@ 1 Introduction
The high latitude (HL) of the Milky Way, also called the “underwater iceberg” guards its secrets about molecular gas and star formation, due in part to the limited scope of previous CO surveys (Xu et al., 2021). Observational challenges essentially originate from the large area of the HL, leading to much longer integration times compared to those required for blind surveys of the Galactic plane. The Planck satellite provides an unprecedented all-sky census of the coldest (6–20 K with a median value of 14 K; Planck Collaboration et al., 2011a) Galactic objects by combining the highest-frequency channels of the Planck survey 353–857 GHz (i.e., 850–350 m) with the far-infrared IRAS 100 m data (Neugebauer et al., 1984; Miville-Deschênes & Lagache, 2005). As a result, the Planck team has cataloged 13,188 Planck Galactic Cold Clumps (PGCCs; Planck Collaboration et al., 2016), including 793 with absolute value of Galactic latitude higher than 30, a group we refer to as high-latitude Planck cold clumps (HLPCs). Benefiting from the unbiased nature of its parent sample, the 793 HLPCs are the least-biased sample of HL cold dust clumps, therefore serving as a foundation for studying the properties of the HL molecular gas and investigating the initial condition of star formation (Wu et al., 2012; Liu et al., 2013).
@@
Our previous work performed a CO/CO/CO (1-0) survey toward 41 early cold cores (ECCs; i.e., most reliable detections of PGCCs with signal-to-noise ratio ; Planck Collaboration et al., 2011b) with the Purple Mountain Observatory (PMO) 13.7 m millimeter-wavelength telescope (Xu et al., 2021). Although detected CO cores have a typical density of several times cm, consistent with what has been found in nearby molecular cloud cores (Benson & Myers, 1983; Myers et al., 1983; Myers & Benson, 1983; Myers, 1983; Benson & Myers, 1989), the turbulent energy is significantly higher than the gravitational energy, with median virial parameters of (Xu et al., 2021). Therefore, our CO surveys unveiled a highly turbulent, diffuse molecular gas environment as a first glimpse of the initial conditions of star formation in the HL clouds.
@@
Stars form in dense cores (Shu et al., 1987) with a typical size of pc and density of – cm (Ward-Thompson et al., 1994, 1999; Kirk et al., 2005). The CO (1-0) transitions suffer from optical thickness and depletion at low temperature, so it is hard to probe the densest regions of molecular clouds. For example, a high-latitude cloud L1780 shows a cometary morphology and a CO core (Toth et al., 1995) but contains no dense core in our survey (field G358.96+36.81). Furthermore, the angular resolution of Planck used for the extraction of PGCCs is ′, corresponding to 0.3 pc at a typical distance of HL clouds of 200 pc (Xu et al., 2021), which is marginal for resolving dense cores. Working at 850 m with an effective beam FWHM of 146, Submillimetre Common-User Bolometer Array 2 (SCUBA-2; Dempsey et al., 2013) provides times better resolution than the Planck, pinpointing cold dense cores inside the molecular clouds.
@@ 2.1 Sample Selection
A thorough search of the SCUBA-2 850 m data in the JCMT Science Archive111https://www.cadc-ccda.hia-iha.nrc-cnrc.gc.ca/en/jcmt/, and crossmatching with 793 HLPCs that satisfy the latitude criterion of , gave 138 observing fields in total. We dismiss six of the fields that have limited integration time or nonstandard scan modes. Different scan patterns include constant velocity daisy patterns (CV Daisy) and rotating curvy pong patterns (Curvy Pong), so the field offsets vary significantly between different patterns. We make sure that the observing fields cover the peak of PGCCs at 353 GHz. We also check superposition or repetition: if two fields cover the same PGCC, we choose the one with the higher sensitivity. After the work flow, a total of 70 SCUBA-2 fields are selected as the sample in this work.
@@ 2.2 Sample Properties
Seventy HLPCs are shown with green crosses, overlaid on the background Planck 353 GHz (850 m) continuum emission in Figure 1. White rectangles outline the CO emission regions defined by Dame et al. (1987). The clump-averaged H column density was calculated assuming a dust-emissivity model at 857 GHz (Planck Collaboration et al., 2016). The distributions of three samples–all the PGCCs, 793 high-latitude PGCCs, and 70 HLPCs–are plotted as gray, blue and orange histograms in Figure 2, respectively. 70 HLPCs has been evenly sampled in the space, ensuring a similar distribution with its parent sample, 793 HL PGCCs. More importantly, the studied sample includes the densest clumps at the high-column-density end ( cm). Considering that the denser clumps should be more likely to produce dense cores, we have covered the complete HLPCs where dense cores could form.
@@ 2.3 Distance Estimation
Distance is always a difficult quantity to estimate in astronomy. Previous studies have estimated the distances of HL molecular clouds to be 100 pc from the velocity dispersion and the scale height of an ensemble of clouds (Blitz et al., 1984; Magnani et al., 1985). The star counting confirms that the HL molecular clouds are indeed nearby objects with upper limit ranges from 125 to 275 pc (Magnani & de Vries, 1986). Using Strömgren photometry, Franco (1989) derive the distances of several HGaL clouds to be 100–230 pc. Both the small and the lack of a double-sine wave signature in the distribution on the (Galactic longitude ) plane demonstrate that HL molecular gas belongs to the local interstellar medium (ISM) and is too close to the Sun for Galactic rotation to modulate the velocities (Magnani et al., 1996).
Recently, the Gaia satellite has provided new photometric measurements toward galactic stars (Gaia Collaboration et al., 2016). Together with 2MASS and Pan-STARRS 1 optical and near-infrared photometry, Gaia DR2 parallaxes can help to infer distances and reddenings of 800 million stars. These stars trace the reddening on a small patch of the sky, along different lines of sight and different distance intervals, allowing us to build a 3D dust-reddening map (Green et al., 2014, 2019). In a given direction, a jump of dust reddening is expected at a distance where there is a dust clump. Distances are estimated by this method and are listed in column (9) of Table A1, with an average value of pc, indicating that HLPCs are mostly local ISM. Adopting Eq. 1 in Xu et al. (2021) and considering that the Sun is 10 pc above the Galactic midplane (Griv et al., 2021), the altitude from the midplane (in units of parsecs) is calculated from where is the distance and is the latitude, and listed in column (10). We note that some fields may contain extragalactic objects, so the distance should be only for foreground Galactic dust.
@@ 3.1 Source Extraction
We adopt the dendrogram algorithm (Rosolowsky et al., 2008) to extract dense structures and then measure their integrated flux, peak flux, size, and position by 2D Gaussian fitting. The details of the algorithm parameter settings and the source extraction procedure are introduced in Appendix B. Within the 70 input fields, we have initially detected a total of 20 sources that belong to 15 SCUBA-2 observing fields. The field names and extracted sources are listed in columns (1)–(2) of Table 1. Central coordinates, standard deviation of the deconvolved major and minor axes, integrated flux, and peak intensity are listed in columns (3)–(7).
@@ 3.2 Source Identification and Dense Core Definition
The CV Daisy mode of observation can produce artifacts (Liu et al., 2018; Eden et al., 2019) that are extracted by the algorithm as false source detections. Therefore, we additionally require that both the peak intensity after being smoothed to the Planck beam and the total flux of the source at 353 GHz, are lower than those of the parent PGCC. This results in two sources, G50.41-35.40 SMM1 and G197.98+33.10 SMM1, being classified as artifacts and excluded from further analyses. We note that the flux given by the PGCC catalog should be from the cold residual map (Planck Collaboration et al., 2016), so the original Planck flux at 353 GHz could be even larger. However, considering the cold nature of dense cores, this should contribute little to the warm component.
We crossmatch the true detections within 1′ using SIMBAD222http://simbad.u-strasbg.fr/simbad/. We find that only one field, G6.04+36.77, pointing toward the molecular cloud L183, contains three resolved (or marginally resolved) sources, which were previously identified as low-mass prestellar cores (Dickens et al., 2000; Pagani et al., 2003). Our other detections are unresolved as point sources, classified as either young stellar object (YSO) or extragalactic object (point sources) – including gravitational lensed galaxy (LeG), BL Lacertae object (BLL), protocluster of galaxies (PClG), active nuclei candidate (AGN), and quasar (QSO). The determined identifiers and references are given in columns (8) and (9) of Table 1.
@@ 4.1 Dense Cores Are Scarce In HLPCs
Having only one detection among the 70 HLPCs highlights the scarcity of dense cores at these latitudes. To further confirm this discrepancy compared to the low-latitude () counterpart, 1235 observing fields from the JCMT Large Project “SCUBA-2 Continuum Observations of Pre-protostellar Evolution” (SCOPE; Liu et al., 2018; Eden et al., 2019) are used as a comparison group because of the following two reasons: (1) twenty-one HLPC fields in this work come from the SCOPE project so that they have been observed with comparable sensitivities; (2) the SCOPE observations serve as a representative sample of PGCCs, with similar distributions in distance, size, and temperature, and with complete column density coverage over cm (Liu et al., 2018).
Considering the beam dilution effects for marginally resolved sources, the column densities are corrected by a beam filling factor where is the Planck beam size at 857 GHz (Planck Collaboration et al., 2016) and is the intrinsic size deconvolved from the beam. For extended sources of which intrinsic size exceeds the beam size, is set to 1, and no correction is performed. As a result, the corrected column density increases by a factor of 1.1 on average and 9 at maximum.
@@ 4.2 What Does “Dense” Mean?
The scarcity of detection does not necessarily indicate the scarcity of dense cores, primarily due to two factors: (1) the limited sensitivity of identifying sources and (2) the absence of large-scale flux in the SCUBA-2 data processing. Therefore, to understand what the scarcity of dense cores means, it is essential to clarify what “dense” means.
Prestellar cores are observed to have flat inner-density gradients that approach beyond a few thousand astronomical units (Ward-Thompson et al., 1994, 1999; Kirk et al., 2005), which can be reproduced by a nonmagnetic and Plummer-like model (Whitworth & Ward-Thompson, 2001) as
@@
Using the combined ammonia data from the Green Bank Telescope and Karl G. Jansky Very Large Array, the temperatures of the three prestellar cores are observed to have a minor decrease toward the center of the core au (Lin et al., 2023). Therefore, we consider a constant temperature profile as K in the following discussion.
@@
In Figure 4, the background color map shows observed peak intensities across the 2D parameter space of flat radius au and central density cm. The prestellar cores have been reported to have au (Ward-Thompson et al., 1999; Kirk et al., 2005), which are delineated by the left green dashed line in Figure 4. The right green dashed line marks 20,000 au, which corresponds to 0.1 pc.
@@
To further demonstrate this upper limit of density, we smooth the images to a resolution of 20″. With better sensitivity, two new cores (on in HLPC G159.41-34.37 and one in G161.87-35.76) are identified by the same algorithm and parameter input, which are called weakly detected cores. They have radii of about 6700 and 7500 au and averaged density of cm and cm, respectively. The two cores are labeled as orange rectangles in Figure 4. We also retrieve Herschel cold cores in L134 (HLPCs G4.13+35.75 and G4.17+36.67 in our survey), MBM12 (G159.21-34.28, G158.51-33.99, G159.14-33.79, G159.23-34.51, G159.41-34.37, G159.66-34.31), L1642 (G210.90-36.55), and LDN1780 (G358.96+36.81, G359.21+36.89) from Montillaud et al. (2015) which are labeled as green triangles. As they all lie in the gray region which is below the sensitivity limit, these cold cores are not dense enough for detection and are consistently below the density limit of cm. As noted by Ward-Thompson et al. (2016), SCUBA-2 selects the densest cores from a population at a given temperature, which makes SCUBA-2 ideal for identifying those cores in Herschel catalogs that are closest to forming stars. So it is of great interest to study how these low-density cores form and whether they can still form stars or are transient objects.
As indicated by the black curves, peak intensity of a synthetic model is always below the corresponding outlined by white curves. The difference reflects the missing large-scale flux, which increases in importance from 0.06 to 0.31 dex with from 2500 to 20,000 au. Therefore, if we do not consider missing flux, the density limit can be even lower, especially for those cores with larger flat radius.